In a survey of women in the United States (ages 20-29), the mean height is 69.9 inches with a standard deviation 3.0 inches.
a. What height represents the 95th percentile?
b. What height represents the first quartile?

Answer by stanbon(75887)

In a survey of women in the United States (ages 20-29), the mean height is 69.9 inches with a standard deviation 3.0 inches.
a. What height represents the 95th percentile?
Find the z-value with a left tail of 95%: z = 1.645
Use x = zs+u
x = 1.645*3+69.9 = 74.835 inches
----------------------------------------------------
b. What height represents the first quartile?
Find the z-value with a left tail of 25%: z = -0.6745
x = -0.6745*3+69.9 = 67.877 inches
==================================
Cheers,
Stan H.